OpenStudy (anonymous):

what is the exact value of (20z)/( radical z^13)

OpenStudy (mertsj):

\[\frac{20z}{\sqrt{z ^{13}}}=\frac{20z}{z ^{6}\sqrt{z}}=\frac{20}{z ^{5}\sqrt{x}}=\frac{20}{z ^{5}\sqrt{z}}\times\frac{\sqrt{z}}{\sqrt{z}}=\frac{20\sqrt{z}}{z ^{6}}\]

OpenStudy (anonymous):

Given, \[20z/\sqrt{z ^{13}}\] First, we should simplify the denominator: z^13 means z*z*z*z*z*z*z*z*z*z*z*z*z To be able to take the square root of a term and "pull" it outside of the radical, we must have values that are squared. We can arrange all of these z's into (z^6)^2*z, or \[(z^6)^2\]. this means we can square the term to get z^6 and pull that out front as a coefficient. Thus, we get 20z/(z^6)*sqrt(z) or \[20z/(z^{6}*\sqrt{z})\] Since we have 20z/z^6, we can reduce a z out so our final answer is \[20/(z ^{5}\sqrt{z})\] HTH!

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Mertsj is much more concise! :)

OpenStudy (mertsj):

yw